OpenStudy (anonymous):

how would i find the missing term of the geometric series a1 = 4, an = 324, r = 3

OpenStudy (campbell_st):

well it looks like you need to find the number of terms n use \[T_{n} = ar^{n-1}\] then \[324 = 4 \times 3^{n -1}\] you need to solve for n divide both sides by 4 \[\frac{324}{4} = 3^{n -1}\] or \[81 = 3^{n -1} \] express 81 as a power of 3 \[3^4 = 3^{n -1}\] equate the powers 4 = n - 1 I'll let you finish the question

OpenStudy (anonymous):

ok thanks